Solve equation (u-3)(u+5)=-12
Answer by Edwin McCravy(20081)


   (u - 3)(u + 5) = -12

The left side is factored, but it doesn't help
us because we don't have 0 on the right side.  So
we must "un-factor" the left side, get 0 on the 
right side, then "re-factor" the new left side:


   (u - 3)(u + 5) = -12

u² + 5u - 3u - 15 = -12

     u° + 2u - 15 = -12

      u² + 2u - 3 = 0 
    
   (u + 3)(u - 1) = 0

u + 3 =  0   u - 1 = 0
    u = -3       u = 1